Goode Perk Company

The Goode Perk Company produces several models of coffee makers. There is little difference in the production time required for the various model coffee machines. The plant is designed to produce 160 coffee machines per eight-hour shift, and there are two shifts per working day.

However, the plant does not operate for the full eight hours: the employees take two 12-minute breaks in each shift, one in the first four hours and one in the second four hours; two hours per week are devoted to cleaning the factory and performing maintenance on the machines; one four-hour period every four weeks is devoted to the meeting of the quality circle. The plant usually produces about 3,500 coffee machines per four-week period. You may ignore holidays in solving this problem. The selling price of the product is $199.95. The variable costs per unit are broken down as follows:

- Labor $60.25

- Raw material 25.70

- Purchased component 21.50

- Variable overhead 27.50

The fixed costs total $503,000 per year.

Utilizing the information in question 4 about the Goode Perk Company, what is the company's revenue at the beak-even volume? Hint: you must calculate break-even volume first.

A $300,000

B $1,547,413

C $799,800

D $1,876,531

E None of the above
Answer #1

Break even volume = Total fixed costs/contribution margin per unit

Contribution margin per unit = selling price per unit - variable cost per unit

=199.95 - 60.25-25.70-21.50-27.50 = $65

Break even volume = 503,000/65 = 7,739 units

Hence, revenue at break even volume = 7,739*199.95

= $1,547,413

Hence, the answer is B